# On quasi-isospectrality of potentials and Riemannian manifolds

### v4

@@ 3 No rigidity of Sturm-Liouville with Dirichlet boundary conditions
An easy computation shows that for $q=0$, $z_{n}^{2}(x,q)=2\sin^{2}{(\pi nx)}$, and it is not difficult to prove that the kernel of $D\Lambda_{0}$ is the space of all functions whose Fourier cosine coefficients vanish. This implies that $D\Lambda_{0}$ is a linear isomorphism when restricted to the space of even functions $\mathcal{E}$, [44, p.53]. Therefore $\Lambda|_{\mathcal{E}}$ is a real analytic isomorphism on a neighbourhood of 0 in $\mathcal{E}$. The latter means that an even function is uniquely determined by its spectrum. There are still isospectral potentials to an even one, but they are not even.
@@
The sequence $\kappa(q)=(\kappa_{n})_{n\geq 1}$ changes sign when $q(x)\mapsto q(1-x)$ (this is a reflexion over $x=1/2$), in contrast to $\Lambda(q)$ which is invariant under this reflection. Let $\Theta$ be a map defined by
@@ 4 Quasi-isospectral potentials, after Bilotta, Morassi, and Turco
In Gesztesy-Telsch [27], the authors explain the single and the double commutation methods for general Sturm-Liouville operators on an interval, a line, and a half line. The single commutation method allows us to introduce an eigenvalue inside a spectral gap, but has the disadvantage that it either introduces singularities or restricts the method to the use of strictly positive eigenfunctions. It goes back to the work of Jacobi [32], Darboux [19], Crum [18], Schmincke [47], and Deift [22]. The double commutation method was introduced by Gel’fand and Levitan in 1955 [25], and it was improved by Gesztesy [26] and Gesztesy and Teschl [27]. More recently, Guliyev [30] investigated the properties of this method applied to isospectral Sturm-Liouville problems with boundary conditions involving rational Herglotz-Nevanlinna functions. The name of the method is related to the fact that it relies on a clever factorization of the operator and the commutation relation of its factors. Below, we will explain the general idea of this method. The approach used by Bilotta et al. has the same background idea. It consists of applying Darboux’s lemma twice. Darboux’s lemma, stated in Lemma 4.1, provides a new potential and a new eigenfunction. However, they both have singularities which are not square integrable on $[0,1]$. Despite that, in the same way as in the second commutation method, Bilotta et al. showed that using twice the lemma, one obtains an expression without singularities, in which all quantities are finite and well defined. Thus, the method leads to regular solutions even though singular functions appear in the process. Let us explain here the parts of the construction that we consider relevant for this paper. The method is also described in [44].
@@ 4.1 Construction of the potential and the eigenfunctions
Let $\mu\in\mathbb{R}$ be fixed, let $g$ be a non trivial solution to
@@
The second step consist in applying the lemma to $z_{m}$ in the place of $f$ and $z_{n}$ in the place of $g$. In that way, we obtain a function $\varphi$ that satisfies
@@
from which is clear that $\kappa_{m,t}(x)$ is well defined and has no singularities in $\mathcal{L}^{2}[0,1]$.
@@
are well-defined and are non trivial solution of the problem
@@ 5.2 Examples with different boundary conditions
Thus $y(0)=c_{1}$, $y^{\prime}(0)=c_{2}\sqrt{\lambda}$ and the first boundary condition gives:
@@
As in previous cases, this problem has a discrete spectrum $(\lambda_{k})_{k\geq 1}$, each eigenvalue being simple, [24]. Let $(\phi_{k})_{k\geq 1}$ be the corresponding eigenfunctions. Applying the lemma of Darboux to $g=\phi_{1}$ and $f=\phi_{k}$, for each $k\geq 2$, we obtain $h_{k}=\frac{1}{\phi_{1}}W[\phi_{1},\phi_{k}]$ is a non trivial solution of
@@ 6.1 Compactness of isospectral Riemannian manifolds
The notion of compactness is often obtained either via Sobolev embeddings and Rellich’s compactness theorem, or through compactness theorems specifically designed for manifolds or metric spaces. One of the best-known results in this area is Gromov’s compactness theorem, which provides conditions -namely, uniform bounds on diameter and on the number of balls of a fixed radius covering the space- under which a sequence of metric spaces admits a convergent subsequence in the Gromov–Hausdorff topology. Other compactness results depend on the particular nature of the problem and the geometric quantities one seeks to control. For example, the Cheeger–Gromov compactness theorem gives sufficient conditions to extract a smoothly convergent subsequence of Riemannian manifolds, assuming uniform bounds on geometric quantities such as Ricci curvature, scalar curvature, diameter, and injectivity radius.
@@ 7 Quasi-isospectral Riemannian manifolds
and denote $\lambda_{\iota}(g)=\lambda_{\iota}$. Therefore
@@
Since we have stablished above that the coefficients corresponding to $j<1+\frac{n}{2}$ vanish, this equation becomes
@@
The equality of the heat invariants follows from the first equation. Since $\lambda_{\iota}>0$, we obtain $s_{1}=s_{2}$. Hence $h_{1}$ and $h_{2}$ are isospectral.
@@
and for $x>x_{0}$, $\varphi(x)$ decreases. Therefore there is a $K>0$ such that for $k\geq K$
@@
see [45]. Isospectral operators have the same determinant, and this is used in the proof of compactness of isospectral metrics on closed surfaces by Osgood et al in [43].
@@
Notice that the series above actually converges for all $s\in\mathbb{C}$. Thus
@@ 8 Quasi-isospectral potentials on Manifolds
are trace class and the corresponding traces have asymptotic expansions for small values of $t$ as given in equation (11). Moreover, using Duhamel’s principle one can also prove that $e^{-t(\Delta+V_{\ell})}-e^{-t(\Delta+V)}$ is trace class and it also has an asymptotic expansion as $t\to 0^{+}$ of the form given in equation (12) since the same proof works in this setting. In the same way as in the proof of Theorem 7.2, we have that
@@ 9.2 Proof of Darboux Lemma
Acknowledgments
The authors acknowledge the support of the “Fundación para la promoción de la investigación y la tecnología” from the Banco de la República de Colombia through the project number 4.498, contract number 202016. We are grateful to the two anonymous referees for pointing out a mistake in the statement of Theorem 7.2, which led to an improvement of the results presented in this paper, and for their constructive comments that helped to enhance its overall quality.

### v5

@@ Abstract
Acknowledgements.
The authors acknowledge the support of the “Fundación para la promoción de la investigación y la tecnología” of the Banco de la República de Colombia through project number 4.498, contract number 202016. We are grateful to the two anonymous referees for pointing out an inaccuracy in the statement of Theorem 7.2, which led to an improvement of the results presented in this paper, and for their constructive comments that helped enhance its overall quality.
This article is the result of work that began in the second semester of 2020, during the Covid-19 pandemic. We would like to thank all those around us whose support helped us through this challenging period. Finally, we are grateful to the editor Mariel Saez, for suggesting that we consider this volume for submission.
@@ 3 No rigidity of Sturm-Liouville with Dirichlet boundary conditions
An easy computation shows that for $q=0$, $z_{n}^{2}(x,q)=2\sin^{2}{(\pi nx)}$, and it is not difficult to prove that the kernel of $D\Lambda_{0}$ is the space of all functions whose Fourier cosine coefficients vanish. This implies that $D\Lambda_{0}$ is a linear isomorphism when restricted to the space of even functions $\mathcal{E}$, [44, p.53]. Therefore, $\Lambda|_{\mathcal{E}}$ is a real analytic isomorphism on a neighbourhood of 0 in $\mathcal{E}$. The latter means that an even function is uniquely determined by its spectrum. There are still isospectral potentials to an even one, but they are not even.
@@
The sequence $\kappa(q)=(\kappa_{n})_{n\geq 1}$ changes sign when $q(x)\mapsto q(1-x)$ (this is a reflection over $x=1/2$), in contrast to $\Lambda(q)$ which is invariant under this reflection. Let $\Theta$ be a map defined by
@@ 4 Quasi-isospectral potentials, after Bilotta, Morassi, and Turco
In Gesztesy-Telsch [27], the authors explain the single and the double commutation methods for general Sturm-Liouville operators on an interval, a line, and a half line. The single commutation method allows us to introduce an eigenvalue inside a spectral gap, but has the disadvantage that it either introduces singularities or restricts the method to the use of strictly positive eigenfunctions. It goes back to the work of Jacobi [32], Darboux [19], Crum [18], Schmincke [47], and Deift [22]. The double commutation method was introduced by Gel’fand and Levitan in 1955 [25], and it was improved by Gesztesy [26] and Gesztesy and Teschl [27]. More recently, Guliyev [30] investigated the properties of this method applied to isospectral Sturm-Liouville problems with boundary conditions involving rational Herglotz-Nevanlinna functions. The name of the method is related to the fact that it relies on a clever factorization of the operator and the commutation relation of its factors. Below, we will explain the general idea of this method. The approach used by Bilotta et al. has the same background idea. It consists of applying Darboux’s lemma twice. Darboux’s lemma, stated in Lemma 4.1, provides a new potential and a new eigenfunction. However, they both have singularities which are not square integrable on $[0,1]$. Despite that, in the same way as in the second commutation method, Bilotta et al. showed that using twice the lemma, one obtains an expression without singularities, in which all quantities are finite and well-defined. Thus, the method leads to regular solutions even though singular functions appear in the process. Let us explain here the parts of the construction that we consider relevant for this paper. The method is also described in [44].
@@ 4.1 Construction of the potential and the eigenfunctions
Let $\mu\in\mathbb{R}$ be fixed, let $g$ be a non-trivial solution to
@@
The second step consists in applying the lemma to $z_{m}$ in the place of $f$ and $z_{n}$ in the place of $g$. In that way, we obtain a function $\varphi$ that satisfies
@@
from which is clear that $\kappa_{m,t}(x)$ is well-defined and has no singularities in $\mathcal{L}^{2}[0,1]$.
@@
are well-defined and are non-trivial solution of the problem
@@ 5.2 Examples with different boundary conditions
Thus, $y(0)=c_{1}$, $y^{\prime}(0)=c_{2}\sqrt{\lambda}$ and the first boundary condition gives:
@@
As in previous cases, this problem has a discrete spectrum $(\lambda_{k})_{k\geq 1}$, each eigenvalue being simple, [24]. Let $(\phi_{k})_{k\geq 1}$ be the corresponding eigenfunctions. Applying the lemma of Darboux to $g=\phi_{1}$ and $f=\phi_{k}$, for each $k\geq 2$, we obtain $h_{k}=\frac{1}{\phi_{1}}W[\phi_{1},\phi_{k}]$ is a non-trivial solution of
@@ 6.1 Compactness of isospectral Riemannian manifolds
The notion of compactness is often obtained either via Sobolev embeddings and Rellich’s compactness theorem, or through compactness theorems specifically designed for manifolds or metric spaces. One of the best-known results in this area is Gromov’s compactness theorem, which provides conditions -namely, uniform bounds on diameter and on the number of balls of a fixed radius covering the space under which a sequence of metric spaces admits a convergent subsequence in the Gromov–Hausdorff topology. Other compactness results depend on the particular nature of the problem and the geometric quantities one seeks to control. For example, the Cheeger–Gromov compactness theorem gives sufficient conditions to extract a smoothly convergent subsequence of Riemannian manifolds, assuming uniform bounds on geometric quantities such as Ricci curvature, scalar curvature, diameter, and injectivity radius.
@@ 7 Quasi-isospectral Riemannian manifolds
and denote $\lambda_{\iota}(g)=\lambda_{\iota}$. Therefore,
@@
Since we have established above that the coefficients corresponding to $j<1+\frac{n}{2}$ vanish, this equation becomes
@@
The equality of the heat invariants follows from the first equation. Since $\lambda_{\iota}>0$, we obtain $s_{1}=s_{2}$. Hence, $h_{1}$ and $h_{2}$ are isospectral.
@@
and for $x>x_{0}$, $\varphi(x)$ decreases. Therefore, there is a $K>0$ such that for $k\geq K$
@@
see [45]. Isospectral operators have the same determinant, and this is used in the proof of compactness of isospectral metrics on closed surfaces by Osgood et al. in [43].
@@
Notice that the series above actually converges for all $s\in\mathbb{C}$. Thus,
@@ 8 Quasi-isospectral potentials on Manifolds
are trace class and the corresponding traces have asymptotic expansions for small values of $t$ as given in equation (11). Moreover, using Duhamel’s principle one can also prove that $e^{-t(\Delta+V_{\ell})}-e^{-t(\Delta+V)}$ is trace class, and it also has an asymptotic expansion as $t\to 0^{+}$ of the form given in equation (12) since the same proof works in this setting. In the same way as in the proof of Theorem 7.2, we have that